OpenStudy (anonymous):

Use the upper and lower sums of 8 rectangles to find two approximations of the area between f(x)=(sqrtx)+2 and the x-axis over the interval [0,2]

OpenStudy (zzr0ck3r):

so for the left sided triangles we have (2-0)/8 = 1/4, so the base of our triangles will be (1/4) and we want to multiply that by our height, which is given by f(x) and since we want the left side, we have (1/4)(f(0)+f(1/4)+f(2/4)+f(3/4)+f(1)+f(5/4)+f(6/4)+f(74/)) = ?

OpenStudy (zzr0ck3r):

replace the word triangle with rectangle....sorry

OpenStudy (zzr0ck3r):

similarly we have for the right \((1/4)(f(1/4)+f(2/4)+f(3/4)+f(1)+f(5/4)+f(6/4)+f(7/4)+f(2)) = ?\)